In a chemistry class, 6 liters of 4% silver iodide solution must be mixed with 10% solution to get a 6% solution. How many liters of 10% solution are needed?
Answer by mananth(16946)

In a chemistry class, 6 liters of 4% silver iodide solution must be mixed with 10% solution to get a 6% solution. How many liters of 10% solution are needed?
Let the quantity of 10% solution be x liters
4% solution------------------- 10% solution---------------- 6% solution ( mix)
6------------------------------- x ------------------------ (x+6) liters

0.04x + 0.1x = 0.06*(x+6)
0.14x= 0.06x+0.36
0.14x-0.06x=0.36
0.08x= 0.36
x= 0.36/0.08
= 4 liters